I need your help desperately.

1. Consider the following equation,
n(7x-3)=21x+5.
For what value of n will the equation result in no solution?
2. Consider the following equation,

n(3x+2)= -9x-6.
For what value of n will the equation result in infinitely many solutions?
Please, help me with these questions. Explain in detail.
Thank you.

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)
Answer by ikleyn(52781)

I need your help desperately.
1. Consider the following equation,
n(7x-3) = 21x+5.
For what value of n will the equation result in no solution?

2. Consider the following equation,
n(3x+2) = -9x-6.
For what value of n will the equation result in infinitely many solutions?
Please, help me with these questions. Explain in detail.
Thank you.
~~~~~~~~~~~~~~~~~~~~

(1)  Equation (1) has no solution if and only if n= 3.


     Indeed, if n= 3, then the equation takes the form

          3*(7x-3) = 21x - 5

     or, equivalently,

          21x - 9 = 21x - 5.


     The last equation, OBVIOUSLY, has no solution;  otherwise, -9 = -5, which is absurd.

Part (1) is complete.

--------------

Part  (2)  is very similar.   Try to solve it on your own.

You may report me your answer for part  (2) - I will check it.


////////////////////////


Comment from student: My work for part 2. n(3x + 2) = -9x - 6 3nx + 2n = -9x - 6

If n = -3, we get the following: -9x - 6 = -9x - 6 So, when n = -3, the equation has an infinite amount of solutions.

Is this right? I need to know tonight. Thanks.



My response.   Your work is correct.  Everything is   RIGHT.

Glad to see your perfect work.


Keep going.


|||||||||||||||||||||||||||||


I also got your another comment, posted before your last comment.

These comments are so different by the tone, that I could not believe
that they are written by the same person.